# Finite simple groups acting with fixity $4$ and their occurrence as groups of automorphisms of Riemann surfaces

### v1

@@ 3 Analysis of series of groups
Theorems 3.7 and 3.8 in [24, pp. 92] tell us that $G$ has a maximal subgroup $M$ of structure $\textrm{PSp}_{2}(q^{2}).C_{2}\cong\textrm{PSL}_{2}(q^{2}).C_{2}$. Then Lemma 3.10 gives elements $x,y,z\in M$ that generate $M$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]z=1$.
@@
We recall that $x,y,z\in M$ generate $M$ with additional properties, and if we take $g\in C_{G}(x)\setminus M$, then $[x,g\cdot y]z=[x,y]z=1$. Therefore the tuple $((x),(g\cdot y),(z))$ satisfies Lemma 3.1 (a) and (b)for the list $l$.
@@
Suppose that $q$ is odd. Then $\alpha=\frac{q^{2}+1}{2}$ and $2\alpha$ divides $\left|T\right|$ as well as $\left|U\right|$. If $q\geq 5$, then $\alpha\geq 13$ and we apply Theorem 3.8 of [24] together with formulae for the orders of classical groups: Then $U\cong M\cong\textrm{PSL}_{2}(q^{2}).C_{2}$ or $q=3$, $\alpha=5$ and $U\cong{C_{2}}^{4}.\Omega_{4}^{-}(2)\cong{C_{2}}^{4}.\textrm{Alt}_{5}$, checking the maximal subgroups from Theorem 3.8 in [24, p. 93] for $\alpha$ dividing their order. Now suppose that $q=3$. We let $C_{1}$ denote a conjugacy class of $G$ of elements of order $\alpha=5$ and we let $C_{2}$ denote a conjugacy class of elements of order $9$. Then we calculate $d_{121}=585$ and $d_{212}=567$, and we see that proper subgroup of $G$ cannot intersect both $C_{1}$ and $C_{2}$ trivially. Then Lemma 3.2 gives that $l$ has a witness.
@@
Moreover $\left|C_{U}(x)\right|=2q^{2}=\left|C_{M}(x)\right|$ and $\left|C_{U\cap M}(x)\right|\leq 4$, since $o(x)=2$ and $\alpha$ is odd. We had already seen earlier that $\left|C_{G}(x)\right|\geq q^{4}$, and then $\left|C_{G}(x)\setminus(U\cup M)\right|\geq 4$. Choosing $g\in C_{G}(x)\setminus M$ as above, such that $g\cdot y\notin U\cup M$, it follows that $T=G$.
@@
By Theorem 3.11 in [24, pp. 94], $G$ has a maximal subgroup $M$ of structure $\textrm{P$\Omega$}_{4}^{-}(q^{2}).C_{2}\cong\textrm{PSL}_{2}(q^{4}).C_{2}$. Then we use Lemma 3.10 and we find $x,y,z\in M$ that generate $M$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]z=1$. If $q$ is even, then the same lemma tells us that $\left|C_{M}(x)\right|=2q^{4}$. If $q$ is odd, then $q^{4}\equiv 1$ modulo $4$ and $\left|C_{M}(x)\right|=2(q^{4}-1)$. Moreover $C_{G}(x)$ has a section isomorphic to $\Omega_{6}^{-}(q)$, to $\Omega_{6}^{+}(q)$ or to $\Omega_{4}^{-}(q)\times\Omega_{4}^{+}(q)$ in this case, as we can see in Table 4.5.1 in [7]. All these sections have order divisible by $q$.
Returning to the case where $q$ is even, we see that $C_{G}(x)$ has a section isomorphic to $\textrm{Sp}_{2}(q)\times\textrm{P$\Omega$}_{4}^{-}(q)$, to $\textrm{Sp}_{4}(q)$ or to $\textrm{Sp}_{2}(q)\cong\textrm{SL}_{2}(q)$, by (8.6), (8.8), (8.10) and (7.6) in [1]. In particular these sections have order divisible by $q^{2}-1$. Since $\left|C_{G}(x)\right|>\left|C_{M}(x)\right|$, we find an element $g\in C_{G}(x)\setminus M$ and then $[x,g\cdot y]z=[x,y]z=1$. The tuple $((x),(g\cdot y),(z))$ satisfies Lemma 3.1(a) and (b) now, and we only need to check (c). Set $T:=\langle x,g\cdot y,z\rangle$ and assume for a contradiction that $T$ is contained in a maximal subgroup $U$ of $G$. We recall that $\langle x,z\rangle\leq M$. Since $y\in M$ and $g\notin M$, it follows that $T\nleq M$.
Now we use that $G$ is a section of $\textrm{GO}_{8}^{-}(q)$ and that we have a list of maximal subgroups in Theorem 3.11 in [24, pp. 94]. Let $\hat{M}$ denote a maximal subgroup of $\textrm{GO}_{8}^{-}(q)$ such that $U$ is a section of $\hat{M}$. Since $T\leq U$, we see that $\left|T\right|=\alpha=d^{-1}(q^{4}+1)$ divides $|\hat{M}|$. Also, we can exclude the cases (v) to (vii), (x) and (xi) from Theorem 3.11. Moreover (i) to (iv) and (viii) cannot occur because, in these cases, $\alpha$ would not divide $|\hat{M}|$. The final remaining case is $\hat{M}\cong\textrm{GO}_{4}^{-}(q^{2}).C_{2}$ and $U\cong\textrm{P$\Omega$}_{4}^{-}(q^{2}).C_{2}\cong\textrm{PSL}_{2}(q^{4}).C_{2}\cong M$.
@@
Assume otherwise and let $M$ be a maximal subgroup of $G$ containing $U$. By our hypothesis in (U), both $\left|U\right|$ and $\left|M\right|$ are divisible by $\alpha\cdot\beta=(q-1)(q-2s+1)$, by $\alpha\cdot\gamma=q^{2}+1$ or by $\beta\cdot\gamma=(q-1)(q+2s+1)$. Then Theorem 4.1 in [24, p. 117] (if $q>8$) and [5, p. 28] (if $q=8$) immediately give a contradiction.
@@
Let $C_{1}$ denote a conjugacy class of elements of order $\alpha$. Theorem 4.3 (xi) in [24, p. 144] gives us an element $x\in G$ of order $\beta:=q^{2}+q+1$ in such a way that $\left|C_{G}(x)\right|$ is divisible by $\beta^{2}$ and that $C_{G}(x)$ lies in a maximal subgroup of $G$ of structure $(C_{\beta}\times C_{\beta}):\textrm{SL}_{2}(3)$. Then we set $C_{2}:=x^{G}$.
First we prove that any subgroup $U$ of $G$ that contains elements of order $\alpha$ and $\beta$ must be $G$. If this were not the case, then $G$ would have a maximal subgroup $M$ of order divisible by $\alpha=q^{4}-q^{2}+1\geq 2^{4}-2^{2}+1=13$, and then the formulae for $\left|\textrm{SL}_{2}(q^{3})\right|$, $\left|\textrm{SL}_{2}(q)\right|$, $\left|G_{2}(q)\right|$, $\left|\textrm{PGL}_{3}^{\pm}(q)\right|$, $\left|\textrm{PSL}_{2}(q^{3})\right|$, $\left|\textrm{PSL}_{2}(q)\right|$, $\left|\textrm{SL}_{3}^{\pm}(q)\right|$, $\left|C_{\beta}\right|$, $\left|C_{q^{2}-q+1}\right|$ and $\left|\textrm{SL}_{2}(3)\right|=8\cdot 3$ leave only one possibility: $M$ must have structure $C_{\alpha}:C_{4}$, by Theorem 4.3 in [24, p. 144]. But $\beta=q^{2}+q+1\geq 2^{2}+2+1=7$ also divides $|M|$, which is impossible now.
@@
We let $C_{1}$ denote a conjugacy class of elements of order $\alpha$. Using Theorem 4.2 (iv) in [24, p. 138] we take $\beta:=q+3s+1$ and we let $C_{2}$ denote a conjugacy class of elements of order $\beta$.
Let $U\leq G$ be such that $U$ contains elements of orders $\alpha$ and $\beta$. Then we prove that $U=G$: Theorem 4.2 in [24, p. 138] shows that every maximal subgroup of $G$ that has an element of order $\beta$ has structure $C_{\beta}:C_{6}$. This is because $\beta$ divides neither $\left|\textrm{PSL}_{2}(q)\right|$ nor $\left|{}^{2}G_{2}(q_{0})\right|$ (if $q_{0}$ is a proper divisor of $q$). Since $q\geq 27>6$ and $U$ also has an element of order $\alpha\geq 13$, it cannot be contained in a maximal subgroup of $G$.
@@ 4 Proof of the main result
In the cases without mixed fixity, there is only one non-regular orbit and hence a single possibility for the branching datum. This gives exactly the lists in Lines 6, 7, 9 –11 and 13 of Table LABEL:DataList.
@@
If $G=\textrm{PSL}_{2}(7)$ and there is only one non-regular orbit, then this gives Line 2a of Table LABEL:DataList. There is also exactly one possibility for a fixity 2 action, with point stabilisers of order $3$. Under the global fixity 4 hypothesis it is possible for $G$ to act on up to two orbits with fixity 2 and with the same point stabilisers, and this gives the lines 2b and 2c. The group can also act with fixity 3 and point stabilisers of order 7, but then this can oly happen on a single orbit. This gives rise to line 2d, with fixity 3 and 4 combined, and if we allow $G$ to act with fixity 2, 3 and 4 on different orbits, then this gives the branching data in Lines 2e and 2f.
In the case where $G=\textrm{PSL}_{2}(8)$, we have a similar situation. A unique non-regular orbit with fixity 4 action gives Line 3a of Table LABEL:DataList. A combination with a fixity 2 action, for which we have two possible point stabilisers, leds to up to five non-regular orbits and hence to Lines 3b–3i.
@@
If $q$ is a prime power, $q\geq 9$, $q$ is odd and $G=\textrm{PSL}_{2}(q)$, then following Table LABEL:GroupList we see two possibilities for a fixity 4 action, and if $G$ has only one non-regular orbit, then this gives Lines 5a and 5b of Table LABEL:DataList. In the mixed fixity cases, we first suppose that $q\equiv 1$ modulo $4$. Then there can be up to two orbits where $G$ acts with fixity 2, and this gives the possibilities in Lines 5c and 5d. Similarly, if $q\equiv-1$ modulo $4$, then we obtain the branching data in Lines 5e and 5f.
The next case is $G=\textrm{Sz}(q)$, where $q$ is a power of $2$ with odd exponent. There are two possibilities for a fixity 4 action, with point stabilisers of different orders, which means that there can be up to two orbits on which $G$ acts with fixity 4. This gives Lines 8a–8c of table LABEL:DataList. It is also possible for $G$ to act with fixity 2, but only with point stabilisers of order $q-1$, and this can only happen on up to two orbits. Combining up to two orbits with fixity 4 and up to two orbits with fixity 2 gives Lines 8d–8i.
There is only one group left that can act with mixed fixity, namely $\textrm{M}_{22}$. A unique non-regular orbit leads to Line 12a of Table LABEL:DataList, and it is also possible to have one orbit with fixity 4 and another one with fixity 3, as in Line 12b. However, it is possible for $G$ to have further non-regular orbits on which it acts with fixity at most 2.
@@
We recall that the genus $g$ is at least 2 in all lists of the table. The GAP code that we use for the individual groups is given in the final section of this article, with some comments.
In each case we explain why the list has a witness, mostly by writing down a generating tuple that satisfies Lemma 3.1 (unless it takes up too much space) or by referring to one of our previous lemmas.
@@
We begin with the case where $G=\textrm{Alt}_{7}$. Then GAP shows that the generating tuple $(((1,2)(3,7,5)(4,6)),((1,3,7,4)(2,6)),(3,4,5,6,7))$
shows that the list $[\textrm{Alt}_{7},g,1\mid[5,1]]$ has a witness.
For the list $[\textrm{Alt}_{7},g,1\mid[5,1],[7,1]]$ we find the generating tuple
$(((1,3,6,4,2)),((2,4)(6,7)),((2,5,4,6,3),(1,2,3,4,5,7,6)))$.
Next we let $G:=\textrm{PSL}_{2}(7)$. Here we have six different lists to look at, and we use GAP again. The generating tuples that show that the lists have witnesses are:
$(((1,2,6,3,5,7,4),(1,8,4)(2,7,3)),((1,4,3,8,2,7,5),(1,5,8)(2,3,6)),\\ ((1,2)(3,4)(5,8)(6,7)))$ for $[\textrm{PSL}_{2}(7),g,2\mid[2,1]]$,
$(((1,7,4)(2,5,8)),((1,7,2,4)(3,6,8,5)),((1,3)(2,8)(4,6)(5,7),(3,5,7)(4,6,8)))$
for $[\textrm{PSL}_{2}(7),g,1\mid[2,1],[3,1]]$,
(((1,7,4)(2,5,8)),((1,4,3)(5,6,8)),((1,6)(2,8)(3,5)(4,7),(1,3,4)(5,8,6),(((1,7,4)(2,5,8)),((1,4,3)(5,6,8)),((1,6)(2,8)(3,5)(4,7),(1,3,4)(5,8,6),( ( ( 1 , 7 , 4 ) ( 2 , 5 , 8 ) ) , ( ( 1 , 4 , 3 ) ( 5 , 6 , 8 ) ) , ( ( 1 , 6 ) ( 2 , 8 ) ( 3 , 5 ) ( 4 , 7 ) , ( 1 , 3 , 4 ) ( 5 , 8 , 6 ) ,
(3,5,7)(4,6,8)))(3,5,7)(4,6,8)))( 3 , 5 , 7 ) ( 4 , 6 , 8 ) ) ) for $[\textrm{PSL}_{2}(7),g,1\mid[2,1],[3,2]]$,
(((1,7)(2,3)(4,6)(5,8)),((1,8,4)(2,7,3)),((1,2)(3,6)(4,5)(7,8),(((1,7)(2,3)(4,6)(5,8)),((1,8,4)(2,7,3)),((1,2)(3,6)(4,5)(7,8),( ( ( 1 , 7 ) ( 2 , 3 ) ( 4 , 6 ) ( 5 , 8 ) ) , ( ( 1 , 8 , 4 ) ( 2 , 7 , 3 ) ) , ( ( 1 , 2 ) ( 3 , 6 ) ( 4 , 5 ) ( 7 , 8 ) ,
(2,4,6,5,8,3,7)))(2,4,6,5,8,3,7)))( 2 , 4 , 6 , 5 , 8 , 3 , 7 ) ) ) for $[\textrm{PSL}_{2}(7),g,1\mid[2,1],[7,1]]$,
$((),(),((1,4)(2,6)(3,7)(5,8),(1,2,4)(6,7,8),(2,4,6,5,8,3,7)))$ for
$[\textrm{PSL}_{2}(7),g,0\mid[2,1],[3,1],[7,1]]$ and
$((),(),((1,4)(2,6)(3,7)(5,8),(1,4,2)(6,8,7),(1,4,2)(6,8,7),(2,4,6,5,8,3,7)))$
for $[\textrm{PSL}_{2}(7),g,0\mid[2,1],[3,2],[7,1]]$.
@@
The next nine lines of the table belong to the group $G=\textrm{PSL}_{2}(8)$, with different lists. Again we use GAP for finding suitable generating tuples that show that the lists have a witness. The tuples are:
(((1,5,8,2,4,6,3,7,9),(1,4,6,7,9,2,3)),((1,2,6,7,5,3,4),(2,6,4,5,7,3,9)),(((1,5,8,2,4,6,3,7,9),(1,4,6,7,9,2,3)),((1,2,6,7,5,3,4),(2,6,4,5,7,3,9)),( ( ( 1 , 5 , 8 , 2 , 4 , 6 , 3 , 7 , 9 ) , ( 1 , 4 , 6 , 7 , 9 , 2 , 3 ) ) , ( ( 1 , 2 , 6 , 7 , 5 , 3 , 4 ) , ( 2 , 6 , 4 , 5 , 7 , 3 , 9 ) ) ,
((2,3)(4,6)(5,9)(7,8)))((2,3)(4,6)(5,9)(7,8)))( ( 2 , 3 ) ( 4 , 6 ) ( 5 , 9 ) ( 7 , 8 ) ) ) for $[\textrm{PSL}_{2}(8),g,2\mid[2,1]]$,
(((1,4,9)(2,5,3)(6,8,7)),((1,5,7,4,6,2,3,9,8)),((1,3)(2,7)(4,9)(6,8),(((1,4,9)(2,5,3)(6,8,7)),((1,5,7,4,6,2,3,9,8)),((1,3)(2,7)(4,9)(6,8),( ( ( 1 , 4 , 9 ) ( 2 , 5 , 3 ) ( 6 , 8 , 7 ) ) , ( ( 1 , 5 , 7 , 4 , 6 , 2 , 3 , 9 , 8 ) ) , ( ( 1 , 3 ) ( 2 , 7 ) ( 4 , 9 ) ( 6 , 8 ) ,
(3,5,7,9,4,6,8)))(3,5,7,9,4,6,8)))( 3 , 5 , 7 , 9 , 4 , 6 , 8 ) ) ) for $[\textrm{PSL}_{2}(8),g,1\mid[2,1],[7,1]]$,
$((),(),((1,5)(2,3)(6,8)(7,9),(1,3,2,8,4,9,5),(3,5,7,9,4,6,8)))$ for
$[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2]]$,
(((1,4,2,9,5,3,6)),((1,6,8,9,4,2,5)),((2,4)(3,6)(5,7)(8,9),(((1,4,2,9,5,3,6)),((1,6,8,9,4,2,5)),((2,4)(3,6)(5,7)(8,9),( ( ( 1 , 4 , 2 , 9 , 5 , 3 , 6 ) ) , ( ( 1 , 6 , 8 , 9 , 4 , 2 , 5 ) ) , ( ( 2 , 4 ) ( 3 , 6 ) ( 5 , 7 ) ( 8 , 9 ) ,
(1,2,3,8,5,7,9,6,4)))(1,2,3,8,5,7,9,6,4)))( 1 , 2 , 3 , 8 , 5 , 7 , 9 , 6 , 4 ) ) ) for $[\textrm{PSL}_{2}(8),g,1\mid[2,1],[9,1]]$,
$((),(),((1,3)(2,7)(4,9)(6,8),(1,2,5,8,9,6,3,4,7),(1,2,3,8,5,7,9,6,4)))$ for
$[\textrm{PSL}_{2}(8),g,0\mid[2,1],[9,2]]$,
$((),(),((1,2)(3,9)(4,8)(5,7),(2,4,3,7,8,6,9),(1,2,3,8,5,7,9,6,4)))$ for
$[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,1],[9,1]]$,
((),(),((1,2)(3,9)(4,8)(5,7),(2,3,8,9,4,7,6),(2,9,6,8,7,3,4),((),(),((1,2)(3,9)(4,8)(5,7),(2,3,8,9,4,7,6),(2,9,6,8,7,3,4),( ( ) , ( ) , ( ( 1 , 2 ) ( 3 , 9 ) ( 4 , 8 ) ( 5 , 7 ) , ( 2 , 3 , 8 , 9 , 4 , 7 , 6 ) , ( 2 , 9 , 6 , 8 , 7 , 3 , 4 ) ,
(1,2,3,8,5,7,9,6,4)))(1,2,3,8,5,7,9,6,4)))( 1 , 2 , 3 , 8 , 5 , 7 , 9 , 6 , 4 ) ) ) for $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2],[9,1]]$,
((),(),((1,2)(3,9)(4,8)(5,7),(2,4,3,7,8,6,9),(1,3,5,9,4,2,8,7,6),((),(),((1,2)(3,9)(4,8)(5,7),(2,4,3,7,8,6,9),(1,3,5,9,4,2,8,7,6),( ( ) , ( ) , ( ( 1 , 2 ) ( 3 , 9 ) ( 4 , 8 ) ( 5 , 7 ) , ( 2 , 4 , 3 , 7 , 8 , 6 , 9 ) , ( 1 , 3 , 5 , 9 , 4 , 2 , 8 , 7 , 6 ) ,
(1,4,6,9,7,5,8,3,2)))(1,4,6,9,7,5,8,3,2)))( 1 , 4 , 6 , 9 , 7 , 5 , 8 , 3 , 2 ) ) ) for $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,1],[9,2]]$ and
((),(),((1,2)(3,9)(4,8)(5,7),(2,3,8,9,4,7,6),(2,9,6,8,7,3,4),((),(),((1,2)(3,9)(4,8)(5,7),(2,3,8,9,4,7,6),(2,9,6,8,7,3,4),( ( ) , ( ) , ( ( 1 , 2 ) ( 3 , 9 ) ( 4 , 8 ) ( 5 , 7 ) , ( 2 , 3 , 8 , 9 , 4 , 7 , 6 ) , ( 2 , 9 , 6 , 8 , 7 , 3 , 4 ) ,
(1,3,5,9,4,2,8,7,6),(1,4,6,9,7,5,8,3,2)))(1,3,5,9,4,2,8,7,6),(1,4,6,9,7,5,8,3,2)))( 1 , 3 , 5 , 9 , 4 , 2 , 8 , 7 , 6 ) , ( 1 , 4 , 6 , 9 , 7 , 5 , 8 , 3 , 2 ) ) ) for $[\textrm{PSL}_{2}(8),g,0\mid[2,1],[7,2],[9,2]]$.
For $G=\textrm{PSU}_{4}(3)$ we only have two lists to check, and we do not display the generating tuples here because they take up so much space. But the code is discussed in the next section.
@@
Next we see a series of generic examples, namely $\textrm{PSL}_{2}(q)$, where $q$ is odd and $q\geq 9$, with different variations of branching data. For the lists 5a and 5b we turn to Lemma 3.4 if $q\notin\{9,11,19\}$, and we use Lemmas 3.5, 3.6 and 3.7 for the cases where $q\in\{9,11,19\}$, respectively. Then, for the mixed fixity cases in the lines 5c–5f, we consider Lemmas 3.5 and 3.8.

### v2

@@ 3 Analysis of series of groups
Theorems 3.7 and 3.8 in [25, pp. 92] tell us that $G$ has a maximal subgroup $M$ of structure $\textrm{PSp}_{2}(q^{2}).C_{2}\cong\textrm{PSL}_{2}(q^{2}).C_{2}$. Then Lemma 3.10 gives elements $x,y,z\in M$ that generate $M$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]z=1$.
@@
We recall that $x,y,z\in M$ generate $M$ with additional properties, and if we take $g\in C_{G}(x)\setminus M$, then $[x,gy]z=[x,y]z=1$. Therefore the tuple $((x),(gy),(z))$ satisfies Lemma 3.1 (a) and (b)for the list $l$.
@@
Suppose that $q$ is odd. Then $\alpha=\frac{q^{2}+1}{2}$ and $2\alpha$ divides $\left|T\right|$ as well as $\left|U\right|$. If $q\geq 5$, then $\alpha\geq 13$ and we apply Theorem 3.8 of [25] together with formulae for the orders of classical groups: Then $U\cong M\cong\textrm{PSL}_{2}(q^{2}).C_{2}$ or $q=3$, $\alpha=5$ and $U\cong{C_{2}}^{4}.\Omega_{4}^{-}(2)\cong{C_{2}}^{4}.\textrm{Alt}_{5}$, checking the maximal subgroups from Theorem 3.8 in [25, p. 93] for $\alpha$ dividing their order. Now suppose that $q=3$. We let $C_{1}$ denote a conjugacy class of $G$ of elements of order $\alpha=5$ and we let $C_{2}$ denote a conjugacy class of elements of order $9$. Then we calculate $d_{121}=585$ and $d_{212}=567$, and we see that proper subgroup of $G$ cannot intersect both $C_{1}$ and $C_{2}$ trivially. Then Lemma 3.2 gives that $l$ has a witness.
@@
Moreover $\left|C_{U}(x)\right|=2q^{2}=\left|C_{M}(x)\right|$ and $\left|C_{U\cap M}(x)\right|\leq 4$, since $o(x)=2$ and $\alpha$ is odd. We had already seen earlier that $\left|C_{G}(x)\right|\geq q^{4}$, and then $\left|C_{G}(x)\setminus(U\cup M)\right|\geq 4$. Choosing $g\in C_{G}(x)\setminus M$ as above, such that $gy\notin U\cup M$, it follows that $T=G$.
@@
By Theorem 3.11 in [25, pp. 94], $G$ has a maximal subgroup $M$ of structure $\textrm{P$\Omega$}_{4}^{-}(q^{2}).C_{2}\cong\textrm{PSL}_{2}(q^{4}).C_{2}$. Then we use Lemma 3.10 and we find $x,y,z\in M$ that generate $M$ and such that $o(x)=2$, $o(z)=\alpha$ and $[x,y]z=1$. If $q$ is even, then the same lemma tells us that $\left|C_{M}(x)\right|=2q^{4}$. If $q$ is odd, then $q^{4}\equiv 1$ modulo $4$ and $\left|C_{M}(x)\right|=2(q^{4}-1)$. Moreover $C_{G}(x)$ has a section isomorphic to $\Omega_{6}^{-}(q)$, to $\Omega_{6}^{+}(q)$ or to $\Omega_{4}^{-}(q)\times\Omega_{4}^{+}(q)$ in this case, as we can see in Table 4.5.1 in [7]. All these sections have order divisible by $q$.
Returning to the case where $q$ is even, we see that $C_{G}(x)$ has a section isomorphic to $\textrm{Sp}_{2}(q)\times\textrm{P$\Omega$}_{4}^{-}(q)$, to $\textrm{Sp}_{4}(q)$ or to $\textrm{Sp}_{2}(q)\cong\textrm{SL}_{2}(q)$, by (8.6), (8.8), (8.10) and (7.6) in [1]. In particular these sections have order divisible by $q^{2}-1$. Since $\left|C_{G}(x)\right|>\left|C_{M}(x)\right|$, we find an element $g\in C_{G}(x)\setminus M$ and then $[x,gy]z=[x,y]z=1$. The tuple $((x),(gy),(z))$ satisfies Lemma 3.1(a) and (b) now, and we only need to check (c). Set $T:=\langle x,gy,z\rangle$ and assume for a contradiction that $T$ is contained in a maximal subgroup $U$ of $G$. We recall that $\langle x,z\rangle\leq M$. Since $y\in M$ and $g\notin M$, it follows that $T\nleq M$.
[…]
@@
Assume otherwise and let $M$ be a maximal subgroup of $G$ containing $U$. By our hypothesis in (U), both $\left|U\right|$ and $\left|M\right|$ are divisible by $\alpha\cdot\beta=(q-1)(q-2s+1)$, by $\alpha\cdot\gamma=q^{2}+1$ or by $\beta\cdot\gamma=(q-1)(q+2s+1)$. Then Theorem 4.1 in [25, p. 117] (if $q>8$) and [5, p. 28] (if $q=8$) immediately give a contradiction.
@@
Let $C_{1}$ denote a conjugacy class of elements of order $\alpha$. Theorem 4.3 (xi) in [25, p. 144] gives us an element $x\in G$ of order $\beta:=q^{2}+q+1$ in such a way that $\left|C_{G}(x)\right|$ is divisible by $\beta^{2}$ and that $C_{G}(x)$ lies in a maximal subgroup of $G$ of structure $(C_{\beta}\times C_{\beta}):\textrm{SL}_{2}(3)$. Then we set $C_{2}:=x^{G}$.
First we prove that any subgroup $U$ of $G$ that contains elements of order $\alpha$ and $\beta$ must be $G$. If this were not the case, then $G$ would have a maximal subgroup $M$ of order divisible by $\alpha=q^{4}-q^{2}+1\geq 2^{4}-2^{2}+1=13$, and then the formulae for $\left|\textrm{SL}_{2}(q^{3})\right|$, $\left|\textrm{SL}_{2}(q)\right|$, $\left|G_{2}(q)\right|$, $\left|\textrm{PGL}_{3}^{\pm}(q)\right|$, $\left|\textrm{PSL}_{2}(q^{3})\right|$, $\left|\textrm{PSL}_{2}(q)\right|$, $\left|\textrm{SL}_{3}^{\pm}(q)\right|$, $\left|C_{\beta}\right|$, $\left|C_{q^{2}-q+1}\right|$ and $\left|\textrm{SL}_{2}(3)\right|=8\cdot 3$ leave only one possibility: $M$ must have structure $C_{\alpha}:C_{4}$, by Theorem 4.3 in [25, p. 144]. But $\beta=q^{2}+q+1\geq 2^{2}+2+1=7$ also divides $|M|$, which is impossible now.
@@
We let $C_{1}$ denote a conjugacy class of elements of order $\alpha$. Using Theorem 4.2 (iv) in [25, p. 138] we take $\beta:=q+3s+1$ and we let $C_{2}$ denote a conjugacy class of elements of order $\beta$.
Let $U\leq G$ be such that $U$ contains elements of orders $\alpha$ and $\beta$. Then we prove that $U=G$: Theorem 4.2 in [25, p. 138] shows that every maximal subgroup of $G$ that has an element of order $\beta$ has structure $C_{\beta}:C_{6}$. This is because $\beta$ divides neither $\left|\textrm{PSL}_{2}(q)\right|$ nor $\left|{}^{2}G_{2}(q_{0})\right|$ (if $q_{0}$ is a proper divisor of $q$). Since $q\geq 27>6$ and $U$ also has an element of order $\alpha\geq 13$, it cannot be contained in a maximal subgroup of $G$.
@@ 4 Proof of the main result
In the cases without mixed fixity, there is only one non-regular orbit and hence a single possibility for the branching datum. This gives exactly the lists in Lines 6, 7, 9-11 and 13 of Table LABEL:DataList.
@@
If $G=\textrm{PSL}_{2}(7)$ and there is only one non-regular orbit, then this gives Line 2a of Table LABEL:DataList. There is also exactly one possibility for a fixity 2 action, with point stabilisers of order $3$. Under the global fixity 4 hypothesis it is possible for $G$ to act on up to two orbits with fixity 2 and with the same point stabilisers, and this gives the lines 2b and 2c. The group can also act with fixity 3 and point stabilisers of order 7, but then this can only happen on a single orbit. This gives rise to line 2d, with fixity 3 and 4 combined, and if we allow $G$ to act with fixity 2, 3 and 4 on different orbits, then this gives the branching data in Lines 2e and 2f.
In the case where $G=\textrm{PSL}_{2}(8)$, we have a similar situation. A unique non-regular orbit with fixity 4 action gives Line 3a of Table LABEL:DataList. A combination with a fixity 2 action, for which we have two possible point stabilisers, leads to up to five non-regular orbits and hence to Lines 3b–3i.
@@
If $q$ is a prime power, $\geq 9, qq$ is odd and $G=\textrm{PSL}_{2}(q)$, then following Table LABEL:GroupList we see two possibilities for a fixity 4 action, and if $G$ has only one non-regular orbit, then this gives Lines 5a and 5b of Table LABEL:DataList. In the mixed fixity cases, we first suppose that $q\equiv 1$ modulo $4$. Then there can be up to two orbits where $G$ acts with fixity 2, and this gives the possibilities in Lines 5c and 5d. Similarly, if $q\equiv-1$ modulo $4$, then we obtain the branching data in Lines 5e and 5f.
The next case is $G=\textrm{Sz}(q)$, where $q$ is a power of $2$ with odd exponent. There are two possibilities for a fixity 4 action, with point stabilisers of different orders, which means that there can be up to two orbits on which $G$ acts with fixity 4. This gives Lines 8a-8c of table LABEL:DataList. It is also possible for $G$ to act with fixity 2, but only with point stabilisers of order q1q-1, and this can only happen on up to two orbits. Combining up to two orbits with fixity 4 and up to two orbits with fixity 2 gives Lines 8d-8i.
There is only one group left that can act with mixed fixity, namely $\textrm{M}_{22}$. A unique non-regular orbit leads to Line 12a of Table LABEL:DataList, and it is also possible to have one orbit with fixity 4 and another one with fixity 3, as in Line 12b.
@@
We recall that the genus $g$ is at least 2 in all lists of the table. Using GAP ([6]) we calculate, for some individual groups and the corresponding lists in the table, a generating tuple that satisfies Lemma 3.1. This is mostly based on information about the conjugacy classes of the groups and the function GeneratingMCOrbits from the MapClass package ([11]). We only write a few of the generating tuples down, but all the code that we use is explained in a supplementary file. In all generic cases, we refer to one of our previous lemmas.
@@
We begin with the case where $G=\textrm{Alt}_{7}$. Then GAP gives us the generating tuple (((1,2)(3,7,5)(4,6)),((1,3,7,4)(2,6)),(3,4,5,6,7))(((1,2)(3,7,5)(4,6)),((1,3,7,4)(2,6)),(3,4,5,6,7)), which satisfies Lemma 3.1, and therefore the list $\textrm{Alt}_{7},g,1\mid[5,1]]$ has a witness.
For the list $\textrm{Alt}_{7},g,1\mid[5,1],[7,1]]$ we find the generating tuple
(((1,3,6,4,2)),((2,4)(6,7)),((2,5,4,6,3),(1,2,3,4,5,7,6)))(((1,3,6,4,2)),((2,4)(6,7)),((2,5,4,6,3),(1,2,3,4,5,7,6))), again using GAP.
Next we let $G:=\textrm{PSL}_{2}(7)$. Here we have six different lists to look at, and in the supplementary file we explain how we find generating tuples for each of them, with GAP. We sometimes deviate slightly from our standard method, depending on the exact entries of the lists, because our generic function based on GeneratingMCOrbits might be very slow.
@@
The next nine lines of the table belong to the group $G=\textrm{PSL}_{2}(8)$, with different lists. A subtle detail needs to be taken into account for List 3c, but otherwise we can calculate suitable generating tuples very easily with GAP.
@@
For $G=\textrm{PSU}_{4}(3)$ we have two lists to check, and then we move to our first series of generic examples, namely $\textrm{PSL}_{2}(q)$. We note that $q$ is odd and $q\geq 9$, with different variations of branching data. For the lists 5a and 5b we turn to Lemma 3.4 if $q\notin\{9,11,19\}$, and we use Lemmas 3.5, 3.6 and 3.7 for the cases where $q\in\{9,11,19\}$, respectively. Then, for the mixed fixity cases in the lines 5c–5f, we consider Lemmas 3.5 and 3.8.